OpenStudy (anonymous):

What property of equality would it be if the equation is x-8=14

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (skullpatrol):

$$x-8=14$$

OpenStudy (skullpatrol):

What do you think the question is asking for?

OpenStudy (anonymous):

It gives me options. It's either addition property of equality, subtraction property of equality, multiplication property of equality or, division property of equity

OpenStudy (skullpatrol):

What property of equality would it be if the equation is x-8=22-8 because 22-8=14 right?

OpenStudy (skullpatrol):

If the same number is subtracted from equal numbers the differences are equal. $$a-c=b-c$$